OpenStudy (sealed777):

On May 26, 1934, a streamlined, stainless steel diesel train called the Zephyr set the world's nonstop long-distance speed record for trains. Its run from Denver to Chicago took 13 hours, 4 minutes, 58 seconds, and was witnessed by more than a million people along the route. The total distance traveled was 1633.8 km. What was its average speed in km/h?

OpenStudy (mathmate):

Average speed = total distance / total time If the unit is to be in km/hr, then distance must be in km, and time converted to decimal hours. To convert hr-min-sec to hours, decimal hours = h+(m+s/60)/60 Example: 4 hours 24 min 18s =4+(24+18/60)/60 =4+(24+0.3)/60 ....... remember PEMDAS =4+24.3/60 =4+0.405 =4.405 hours